OpenStudy (anonymous):

A 75 kg man is standing on a scale in an elevator. The elevator begins to accelerate upward at a rate of 3.2 m/s^2. What is the weight of the man during the acceleration?

OpenStudy (anonymous):

N-mg = ma N is the normal force which is the force the scale registers...

OpenStudy (anonymous):

\[R=m(g+a)\] \[75(9.8+3.2) = 75 * 13 = 975N\]